OpenStudy (anonymous):

c(x+2)-5=b(x-3) solve for x. distribute first, right? cx+2c-5+bx-3b. Am I right so far?

OpenStudy (ivettef365):

you are right but you put a (+) instead of = cx+2c-5 = bx-3b

OpenStudy (anonymous):

oops. and then cx-bx=-3b-2c+5?

OpenStudy (ivettef365):

now you need to isolate x, so keep all x's on one side as follows: cx -bx = -3b -2c + 5 yes that is correct

OpenStudy (ivettef365):

now distribute the cx - bx as x(c-b)

OpenStudy (anonymous):

x=-3b-2c+5/c-b?

OpenStudy (ivettef365):

CORRECT

OpenStudy (anonymous):

yay! Thank you!

OpenStudy (ivettef365):

yw :)